a full wave rectifier uses load resistor of 1500 ohm assume the diodes have Rf =10ohm Rr =infinite . the voltage applied to diode is 30V with a fequency of 50 Hz. calculate the AC input


1
Expert's answer
2021-09-17T09:23:56-0400

The RMS current is


"I_{rms}=\frac{V_{max}}{(R_L+R_F)\sqrt2}=0.014\text{ A}."

The AC power input is


"P_i=I_{rms}^2(R_F+R_L)=0.298\text{ W}."